OpenStudy (anonymous):

What is 90 feet per hour converted into yards per day? A. 12.5 yards per day B. 11.25 yards per day C. 720 yards per day D. 2160 yards per day

OpenStudy (anonymous):

Forget it. Its 720 which is c

OpenStudy (anonymous):

you need to figure out the conversion factors, for example \[24 km/h \rightarrow m/s \] \[24\frac{ km }{ h } * \frac{ 1000m }{ 1km } = \frac{ 24000m }{ h } \] the km units cancel out \[\frac{ 2400m }{ h } * \frac{ 1h }{ 3600s} = 0.667\frac{ m }{ s }\]